Use the following information for Taco Swell, Inc., (assume the tax rate is 21 percent) 2017 2018 $17,049 $ 18,798 2,386 2,494 Sales Depreciation Cost of goods sold 5,740 6,741 Other expenses Interest Cash Accounts receivable 11,498 13,512 Short-term notes 1,350 1183 1,115 1,330 8,681 9,277 ook rint 1,684 1,651 payable ences Long-term debt Net fixed assets Accounts payable Inventory Dividends 29,090 35,254 72,792 77,640 6,670 20,441 21,872 2,029 2,324 6,275 For 2018, calculate the cash flow from assets, cash flow to creditors, and cash flow to stockholders. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) Cash flow from assets Cash flow to creditors Cash flow to stockholders 10 INext >
Answer #1

Cash flow from Assets = OCF - Capital Spending - Additions to NWC

= (EBIT + Depreciation - Taxes) - ( Ending Net Fixed Assets- Beginning Net Fixed Assets+ Depreciation) - ( Ending NWC - Beginning NWC)

= (18798-6741-1183-2494+2494-0)-(77640-72792+2494)-(9277+13512+21872-1651-6670-(8681+11498+20441-6275-1684))

= 10874-7342-3679

= -147.00

Cash flow to Creditors = Interest - Net new Borrowings

= Interest - (Ending long term Debt - Beginning long term Debt)

= 1330 - (35254-29090)

= 1330-6164

= -4834.00

Cash flow from Assets = Cash flow to Creditors + Cash flow to Stockholders

-147 = -4834 + Cash flow to Stockholders

Cash flow to Stockholders = -147+4834

= 4687.00